# CityBench: Evaluating the Capabilities of Large Language Models for Urban Tasks

### v2

@@ 4 Benchmark and Experiments
Oceania Sydney 1935 5087 21390 141997 438763 54170
4.1 Settings
Model Deployment To facilitate usage of CityBench, we have implemented local deployment support for the majority of LLMs and VLMs using VLMEvalKit (Duan et al., 2024) and vLLM (Kwon et al., 2023). Additionally, we also support evaluation through the APIs of commercial closed-source models, e.g., OpenAI and commercial open-source models, e.g. DeepInfra §§§https://deepinfra.com/.
Baselines We select well-known LLMs and VLMs as baselines. For VLMs, we select LLaVa-NeXT (Liu et al., 2024c), CogVLM-v2 (Wang et al., 2023a), MiniCPM-LLama3-V-2.5 (OpenBMB, 2024), Qwen-VL-plus and GPT4o. For LLMs, we select LLama3-8B, LLama3-70B, Mistral-7B-v0.2 (Jiang et al., 2023), Mixtral-8x22B-v0.1 (Jiang et al., 2024), DeepSeekv2 (Shao et al., 2024), GPT3.5, and GPT4 (Achiam et al., 2023). We also select representative baselines, including GeoCLIP (Vivanco Cepeda et al., 2024) for street view image geolocalization, RSVA (Wang et al., 2022) for infrastructure inference, RemoteCLIP (Yeh et al., 2020; Liu et al., 2024b) for population prediction, LSTPM (Sun et al., 2020) for mobility prediction and MaxPressure (Varaiya, 2013) for traffic signal control task.
@@ 4.1 Settings
Table 2: Performance of 16 widely-used VLMs on four urban visual tasks in CityBench, where ‘City Infer’ and ‘Loc Infer’ denote the city inference task and the geo-coordinates inference task of street view image, ‘Population’ denotes the geospatial prediction task, ‘Infra’ denotes the infrastructure inference task, ‘Navigation’ denotes the outdoor visual-language navigation task.
@@
City Infer Loc Infer Population Infra Navigation
@@
4.2 Experimental Results
Performance of VLMs on Urban Visual Tasks Main results on urban visual tasks are presented in Table 2. We find three interesting observations: 1) VLMs perform much better than the traditional baselines on the urban visual task, e.g, city inference via street view image and infrastructure inference, which require commonsense and semantic understanding ability of models. 2) But, VLMs perform poorly on tasks requiring precise numerical values, or have a diminished advantage, e.g., the geospatial prediction task on population and precise geo-coordindates prediction task. 3) Most of the VLMs’ performance on all the tasks are not robust, e.g., some VLMs’ performance is close to zero and large model within the same model series not always perform better than the smaller models, e.g., InternVL2-26B performs worse than InternVL2-8B. Finally, we find that the performance variability across different LLMs and VLMs is primarily influenced by the capabilities of the LLMs backbone. For example, in widely used VLMs for urban tasks, Intern2.5-7B consistently outperforms Qwen2-7B and Mistral-7B in most tasks, which can be attributed to Intern2.5-7B’s superior performance in general NLP tasks at the same parameter scale. Similarly, LLaVA-NeXT-8B demonstrates performance comparable to MiniCPM-V2.5-8B, as both models leverage the same LLM backbone, LLaMA3-8B.
Performance of LLMs on Urban Tasks without Visual Input Results on urban task without visual input are presented in Table 3. Compared with the simple urban visual tasks, these urban tasks require more reasoning abilities and professional knowledge of LLMs. From Table 3, we find that LLMs perform much better than the traditional methods in mobility prediction task which is consistent with the results from Wang et al. (Wang et al., 2023b). Besides, different from the results from Lai et al. (Lai et al., 2023), the performance of LLMs on the traffic signal control task seems poor which are far worse than the classic Max-Pressure method. It is noted that our testing scene is the more challenging multi-intersection signal control problem. Finally, as for the GeoQA task for city elements, we find that even the best LLM GPT4-Turbo still has a long way from the performance ceiling. Additionally, we present results from several VLMs for these tasks at the bottom of Table 3. Compared to their original LLM backbones, the performance of VLMs shows some decline. For instance, CogVLM2-19B performs worse than Llama3-8B in both mobility prediction and urban exploration tasks. Therefore, maintaining the general capabilities of LLMs during the training of VLMs should be a key direction to ensure their effectiveness across a wide range of tasks.
Figure 5: Detailed results of LLMs on mobility prediction task(left) for different cities and image geolocalization task(right) for different models.
Error Analysis of LLMs We find that LLMs often display errors such as logic error, format error, invalid action, refusal to answer, and hallucinations. As shown in Figure 6, Llama3-8B exhibited a logical error in its judgment of time in the task mobility prediction. For the urban exploration task, Qwen2-7B refused to choose the option and instead demanded the user to use a navigation service to solve the problem. Intern2.5-7B directly stated that it lacks expertise in this area and needs more information to answer the question. Llama3-8B provided an invalid option in the traffic signal task, rendering CitySimu unable to perform next action. We notice that the most frequent error is Misformatted, and several instances close to 0 in Table 3 are mostly caused by formatting errors. Thus, one of the promising direction is to reduce these error from LLMs to improve their practicality.
Error Analysis of VLMs Urban visual tasks require the model to make decisions directly without going through an explanation process. As a result, as shown in Figure 7, common errors include format errors, invalid actions, refusal to answer and hallucinations. In the image geolocalization task, InternVL2-2B provides a response that do not follow the required format, while Yi-VL-34B gives an irrelevant invalid response. CogVLM2-19B and Yi-VL-34B, in the geospatial prediction and infrastructure inference tasks respectively, repeat the examples provided in the question and refuse to answer the actual question. Due to the response format requirements of the tasks, the most common error made by VLMs in urban visual tasks is misformatted responses.
Geospatial Bias of LLMs and VLMs To further investigate the difference between LLMs and VLMs, we report the detailed results of mobility prediction task and image geolocalization tasks from 13 cities in Figure 5. Based on the above results, we have made several interesting discoveries. First, we find that the performance of different LLMs varies a lot across different cities, no LLM can always perform best in mobility prediction tasks. Second, we find that the performance of VLMs on visual task like image geolocalization task are significantly biased. Most VLMs perform well in major international cities, but poorly in some lesser-known cities (e.g., CapeTown and Nairobi). We provide simple evidence for this phenomenon by analyzing the number of public websites available on Google and Wikipedia. For more details, see Section A.3 in the appendix. The variability in evaluation results demonstrate the necessity of establishing a global evaluation benchmark, and also highlights the potential shortcomings and areas for improvement of LLMs.
@@ 4.2 Experimental Results
Table 3: Performance of LLMs and VLMs on four urban tasks without visual input in CityBench, where ‘Top 1 Acc’ denotes the ‘Top1 Accuracy’ metric, ‘Queue’ denotes the ‘Queue Length’ metric.
Tasks Understanding Planning & Decision-making
GeoQA Mobility Prediction Urban Exploration Traffic Signal
@@
Figure 6: Error case analysis in mobility prediction, urban exploration and traffic signal control tasks.
Figure 7: Error analysis in image geolocalization, geospatial prediction and infrastructure inference tasks.
5 Conclusion
@@ Appendix A Appendix
A.1 Few-Shot Performance of LLMs in CityBench
@@
Table 4: Few-shot performance of several representative LLMs in Beijing.
@@
A.2 Details of Quality Control
@@
A.3 Additional Results of Geospatial Bias Analysis
@@
Table 5: The performance of street view image localization tasks in different cities and its relationship with the amount of data in the ‘training corpus’ of LLMs, where the amount of data in the training corpus is approximated by the number of Google search entries and Wikipedia entries for each city.
@@
A.4 Map building tool in CityData
The map building tool enhances open-source map data to support subsequent behavior simulations, encompassing lane topology recovery, relationship recognition, intersection reconstruction, area of interest (AOI) mapping, point of interest (POI) clustering, basic traffic rule generation, and right-of-way construction. The framework of map building tool is presented in Figure 8.
Figure 8: Framework of map building tool.
Table 6: Detailed information of 8 evaluation tasks in CityBench, including data modality, metric and data instances. Task settings across different cities keep consistent.
@@ A.4 Map building tool in CityData
A.5 Discussion

### v3

@@ 3. Benchmark and Experiments
Sydney 1935 5087 21390 141997 438763 54170
3.1. Settings
Model Deployment To facilitate usage of CityBench, we have implemented local deployment support for the majority of LLMs and VLMs using VLMEvalKit (Duan et al., 2024) and vLLM (Kwon et al., 2023). Additionally, we also support evaluation through the APIs of proprietary models, e.g., OpenAI and open-source models, e.g. DeepInfra  [^7] https://deepinfra.com/ and Siliconflow  [^8] https://siliconflow.cn/.
Baselines We select well-known LLMs and VLMs as baselines. For VLMs, we select LLaVa-NeXT (Liu et al., 2024b), CogVLM-v2 (Wang et al., 2023b), MiniCPM-LLama3-V-2.5 (OpenBMB, 2024), Qwen-VL-plus and GPT4o. For LLMs, we select LLama3-8B, LLama3-70B, Mistral-7B-v0.2 (Jiang et al., 2023), Mixtral-8x22B-v0.1 (Jiang et al., 2024), DeepSeekv2 (Shao et al., 2024), GPT3.5, and GPT4 (Achiam et al., 2023). We also select representative baselines, including GeoCLIP (Vivanco Cepeda et al., 2024) for street view image geolocalization, RSVA (Wang et al., 2022) for infrastructure inference, RemoteCLIP (Yeh et al., 2020; Liu et al., 2024a) for population prediction, LSTPM (Sun et al., 2020) for mobility prediction and MaxPressure (Varaiya, 2013) for traffic signal control task.
@@ 3.1. Settings
Table 2. Performance of 16 widely-used VLMs on four urban visual tasks in CityBench. Here, ‘City’ and ‘Loc.’ represent the city name inference task and the geo-coordinates inference task for street view images, respectively; ‘Population’ refers to the geospatial prediction task; ‘Infra’ denotes the infrastructure inference task; and ‘Navigation’ indicates the outdoor visual-language navigation task. ‘Succ.’ stands for the success rate metric, while ‘Dist.’ represents the distance metric.
@@
City Loc. Population Infra Navigation
@@
Table 3. Performance of LLMs and VLMs on four urban tasks without visual input in CityBench. Here, ‘Top 1’ represents the Top-1 Accuracy metric, ‘Succ.’ denotes the success rate metric, ‘Q’ refers to the Queue Length metric, and ‘TP’ indicates the throughput metric. ‘Und.’ stands for the understanding task.
Tasks Und. Planning & Decision-making
GeoQA Mobility Exploration Traffic Signal
@@
3.2. Overall Performance on CityBench
3.2.1. CityBench are Challenging for LLMs and VLMs
The performance of LLMs and VLMs on CityBench is summarized in Table 3 for urban tasks without visual input and in Table 2 for urban visual tasks. As we can observe, except for the street view image geolocalization and infrastructure inference tasks, the performance of LLMs on the remaining six tasks is suboptimal and remains far from the ideal ceiling. For instance, the performance on GeoQA task, which evaluates detailed knowledge about urban elements of LLMs, is only 0.398, significantly lower than the best possible score of 1.0. These results demonstrate that CityBench is a challenging benchmark on urban tasks for LLMs.
3.2.2. LLMs and VLMs Struggle with Numerical Tasks
As shown in Table 3 and Table 2, the performance of LLMs and VLMs on numerical tasks, including population estimation and traffic signal control, significantly lags behind existing baselines. In population estimation tasks, the best-performing LLM, GPT-4o, underperforms RemoteCLIP by 18% in terms of RMSE. Similarly, in traffic signal control tasks, the top-performing VLM, InternVL2-40B, trails behind the Max-Pressure method by 41.9% in queue length. Therefore, improving the performance of LLMs on numerical tasks is crucial for their application in urban tasks. Although some studies have explored potential solutions by adding task-specific heads to LLMs, such designs may compromise the generalizability of the model.
3.2.3. Performance Consistency between Various Urban Tasks
As the best-performing VLM in urban visual tasks, GPT-4o only excels in the first two tasks, while other VLMs, such as GLM-4v-9B, LLaVANeXT-34B, and InternVL2-8B, outperform it in the remaining two tasks. For urban tasks without visual input, LLama3-70B achieves the best performance in mobility prediction and urban exploration tasks, but it falls behind other high-performing LLMs like GPT-4 Turbo and InternVL2-40B in the other two tasks. In other words, due to the heterogeneity and complexity of urban tasks, no LLM or VLM, including the powerful GPT-4 series, can consistently perform well across all tasks. These results highlight the challenges posed by CityBench and underscore the necessity of developing domain-specific LLMs and VLMs tailored for urban tasks.
3.2.4. LLMs and VLMs Exhibit Geospatial Bias
To further investigate the difference between LLMs and VLMs, we report the detailed results of mobility prediction task and image geolocalization tasks from 13 cities in Figure 5. Based on the above results, we have made several interesting discoveries. First, we find that the performance of different LLMs varies a lot across different cities, no LLM can always perform best in mobility prediction tasks. Second, we find that the performance of VLMs on visual task like image geolocalization task are significantly biased. Most VLMs perform well in major international cities, but poorly in some lesser-known cities (e.g., CapeTown and Nairobi). We provide preliminary evidence for this phenomenon by analyzing the number of publicly accessible websites indexed on Google and Wikipedia. We find that cities with fewer publicly available websites tend to pose greater challenges for LLMs, while those with more extensive online presence are generally easier for the models to handle. For more details, see section A.3 in the appendix. The variability in evaluation results demonstrate the necessity of establishing a global evaluation benchmark, and also highlights the potential shortcomings and areas for improvement of LLMs.
Figure 5. Detailed performance results of LLMs on two tasks: (top) mobility prediction and (bottom) image geolocalization. Both tasks are evaluated across multiple cities and multiple models, demonstrating that significant performance variations across diverse urban contexts are consistently observed even with different model architectures, highlighting the pervasive nature of geospatial bias in these models.
3.3. Detailed Analysis between Models
3.3.1. Disparity between Proprietary and Open-source Models
Although the GPT-4 series performs well on most tasks, several open-source models have achieved better results on multiple other tasks, albeit not consistently by a single open-source model. In other words, we do not observe a dominant advantage of proprietary models over open-source models on CityBench, which may be attributed to the uniqueness and heterogeneity of urban tasks. The reasons behind this phenomenon warrant further in-depth analysis in future research. While the performance of models within the same series generally follows the scaling law with respect to model size, most VLMs’ performance across tasks is not robust. For instance, some VLMs exhibit near-zero performance, and larger models within the same series do not always outperform smaller ones. For example, InternVL2-26B performs worse than InternVL2-8B. Similar trends are observed in LLMs, where Intern2.5-20B does not consistently outperform Intern2.5-7B.
3.3.2. Performance Correlation between VLMs and LLMs
While most VLMs are continuously trained based on LLMs, we investigate the impact of LLMs on VLM performance. First, we find that the performance variability across different LLMs and VLMs is primarily influenced by the capabilities of the LLM backbone. For instance, in widely used VLMs for urban tasks, Intern2.5-7B consistently outperforms Qwen2-7B and Mistral-7B across most tasks, which can be attributed to Intern2.5-7B’s superior performance in general NLP tasks at the same parameter scale. Similarly, LLaVA-NeXT-8B demonstrates performance comparable to MiniCPM-V2.5-8B, as both models share the same LLM backbone, LLaMA3-8B. Second, as shown in Table 3, we observe that most VLMs underperform compared to their LLM bases on urban tasks without visual input. This demonstrates that while post-training of LLM-based VLMs enhances visual capabilities, it inevitably leads to performance degradation in original textual tasks. For example, LLaVA-NeXT-8B lags behind LLaMA3-8B on all textual tasks, with an average performance degradation of 4.10%, while MiniCPMv2.5-8B exhibits a smaller degradation of over 1.88%. Therefore, maintaining the general capabilities of LLMs during the training of VLMs should be a key direction to ensure their effectiveness across a wide range of tasks.
Figure 6. Error analysis in mobility prediction, urban exploration, and traffic signal control tasks reveals common issues: logic errors, format errors, invalid actions, refusal to answer, and hallucinations. Full prompts for each task are in the appendix.
3.3.3. Typical Errors of LLMs and VLMs in CityBench
We find that LLMs often display errors such as logic error, format error, invalid action, refusal to answer, and hallucinations. The types of errors are highly correlated with the characteristics of the LLMs. For instance, certain models, such as MiniCPM2.5-8B, exhibit excessive alignment in handling geospatial-related content, leading to a systematic refusal to respond to queries across various tasks, as illustrated in the image localization task depicted in Figure 5. On the other hand, smaller models like InternVL2-2B often struggle to follow instructions, leading to format errors and invalid actions. We present typical error cases in Figure 6. For example, Llama3-8B exhibited a logical error in its judgment of time in the task mobility prediction. For the urban exploration task, Qwen2-7B refused to choose the option and instead demanded the user to use a navigation service to solve the problem. Intern2.5-7B directly stated that it lacks expertise in this area and needs more information to answer the question. Llama3-8B provided an invalid option in the traffic signal task, rendering CitySimu unable to perform next action. We notice that the most frequent error is Misformatted, and several instances close to 0 in Table 3 are mostly caused by formatting errors. Thus, one of the promising direction is to reduce these error from LLMs to improve their practicality. More detailed analysis on VLMs are presented in section A.4 of appendix.
4. Related Work
Evaluating LLMs for Urban Knowledge and Tasks. Researchers from various urban related fields have conducted extensive evaluations of LLM in urban space from different aspects (Feng et al., 2025b; Ding et al., 2024). Kuckreja et al. (Kuckreja et al., 2023) evaluate the performance of multi-modal LLMs on several remote sensing related tasks. Yang et al. (Yang et al., 2024) propose V-IRL benchmark to evaluate the performance of multi-modal LLMs on street view image related tasks including localization and recognition tasks. Mai et al. (Mai et al., 2023) and Manvi et al. (Manvi et al., 2023) use LLMs to predict social indicators like population and education level. Gurnee et al. (Gurnee and Tegmark, 2023) and Bhandari et al. (Bhandari et al., 2023) try to testify whether LLMs know the coordinates of geospatial entity. Mooney et al. (Mooney et al., 2023) and Deng et al. (Deng et al., 2023) use GIS exams to understand the geospatial skills of LLMs. Different from these works, we first introduce the interactive simulator based systematic evaluation system for LLMs and VLMs, which covers various data modalities, diverse urban task types and differentiated data from 13 cities around the world.
Interactive Decision-making and Urban Simulator. Beyond the above static evaluation, researchers also evaluate the capacity of LLMs in the interactive decision making tasks with customized simulators, e.g., web agent (Liu et al., 2023b) with web environment and embodied intelligence (Yang et al., 2023) with virtual home (Puig et al., 2018). In the urban domain, Schumann et al. (Schumann et al., 2024) apply LLM to do the visual language navigation task in Touchdown (Chen et al., 2019) and Lai (Lai et al., 2023) apply LLMs as the traffic light controller in CityFlow (Zhang et al., 2019) to manage the road traffic. Besides, Yang et al. (Yang et al., 2024) design V-IRL as the environment of street view image related tasks and propose a global scale virtual intelligence benchmark. These works only evaluate the potential of LLMs in single urban decision-making task and most of their results rely on small-scale datasets in limited regions. Different from them, our work builds on an efficient urban simulator with global scale and supports 4 representative urban decision-making tasks with different modality in one benchmark, including urban exploration, outdoor navigation, mobility prediction and traffic control task.
5. Conclusion
@@ Appendix A Appendix
A.1. Few-Shot Performance of LLMs in CityBench
@@
Table 4. Few-shot performance of several representative LLMs in Beijing.
@@
A.2. Details of Quality Control
@@
A.3. Additional Results of Geospatial Bias Analysis
@@
Table 5. The performance of street view image localization tasks in different cities and its relationship with the amount of data in the ‘training corpus’ of LLMs, where the amount of data in the training corpus is approximated by the number of Google search entries and Wikipedia entries for each city.
@@
A.4. Error Analysis of VLMs
Urban visual tasks require the model to make decisions directly without going through an explanation process. As a result, as shown in Figure 7, common errors include format errors, invalid actions, refusal to answer and hallucinations. In the image geolocalization task, InternVL2-2B provides a response that do not follow the required format, while Yi-VL-34B gives an irrelevant invalid response. CogVLM2-19B and Yi-VL-34B, in the geospatial prediction and infrastructure inference tasks respectively, repeat the examples provided in the question and refuse to answer the actual question. Due to the response format requirements of the tasks, the most common error made by VLMs in urban visual tasks is misformatted responses.
Figure 7. Error analysis in image geolocalization, geospatial prediction and infrastructure inference tasks.
A.5. Map building tool in CityData
The map building tool  [^9] https://github.com/tsinghua-fib-lab/mosstool enhances open-source map data to support subsequent behavior simulations, encompassing lane topology recovery, relationship recognition, intersection reconstruction, area of interest (AOI) mapping, point of interest (POI) clustering, basic traffic rule generation, and right-of-way construction.
Table 6. Detailed information of 8 evaluation tasks in CityBench, including data modality, metric and data instances. Task settings across different cities keep consistent.
@@ A.5. Map building tool in CityData
A.6. Discussion
We discuss some limitation of current work as below.